The final answer should be in the format of "tolerance of ± in the third significant digit."


1.The drag racer P starts from rest at the start line S and then accelerates along the track. When it has traveled 87 m, its speed is 39 m/s and it is accelerating forward at 10.1 m/s2. For that instant, determine the values of r¨ and θ¨ relative to axes fixed to an observer O in the grandstand G as shown.


28 meters from point O ( it is the base of the triangle)


 r¨ = ____ m/s2

 θ¨ = _____ rad/s2





1
Expert's answer
2021-04-01T18:34:39-0400

"\theta=atan\frac{SO}{SP}=17.84°,"

"r=OP=\sqrt{OS^2+SP^2}=91.3947~m,"

"\dot{r}=v_r=v cos \theta=37.1247~\frac ms,"

"v_{\theta}=-v sin \theta=-11.9480~\frac ms,"

"v_{\theta}=r\dot{\theta}\implies \dot{\theta}=\frac{v_{\theta}}{r}=-0.1307~\frac ms,"

"a_r=acos \theta=9.6143~\frac{m}{s^2},"

"a_{\theta}=-asin \theta=-3.0942~\frac{m}{s^2},"


"a_r=\ddot{r}-r\dot{\theta}^2,\implies \ddot{r}=a_r+r\dot{\theta}^2=11.176~\frac{m}{s^2},"

"a_{\theta}=r\ddot{\theta}+2\dot{r}\dot{\theta}\implies \ddot{\theta}=\frac{a_{\theta}-2\dot{r}\dot{\theta}}{r}=-0.0723~\frac{rad}{s^2}."